# GUI Agents: A Survey

### v1

@@ 4.3.2 Planning with External Knowledge
Enabling LLM-powered agents to interact with diverse applications and resources through GUIs allows them to leverage external data sources, thereby enhancing their planning capabilities. For example, Search-Agent (Koh et al., 2024b) combines LLM inference with A* search to explore and backtrack to alternative paths explicitly, AgentQ (Putta et al., 2024) combines LLM with MCTS. Toolchain (Zhuang et al., 2023) models tool planning as a tree search algorithm and incorporates A* search to adaptively retrieve the most promising tool for subsequent use based on accumulated and anticipated costs. SGC (Wu et al., 2024a) decomposes the query and performs embedding similarity match between the concatenated subquery with the current retrieved task API and each of the existing APIs, and then selects the top one from the existing neighboring APIs. Thought Propagation Retrieval (Yu et al., 2023) prompts LLMs to propose a set of analogous problems and then applies established prompting techniques, like Chain-of-Thought, to derive solutions. The aggregation module subsequently consolidates solutions from these analogous problems, enhancing the problem-solving process for the original input. WebShop, Mind2Web, and WebArena Zhou et al. (2023c); Deng et al. (2023) allow agents to interact with webs to plan for web browsing for search. WMA (Chae et al., 2024) utilizes world models to address the mistakes made by LLMs for long-horizon tasks.
@@ 4.4 Acting
Acting in GUI agents involves translating the agent’s reasoning and planning outputs into executable steps within the GUI environment. Unlike purely text-based or API-driven agents, GUI agents must articulate their actions at a finer granularity—often down to pixel-level coordinates—while also handling higher-level semantic actions such as typing text, scrolling, or clicking on specific elements. Several directions of approaches have emerged:
@@ 5 GUI Agent Training Methods
This section summarizes different strategies to elicit the ability to solve agentic tasks in a GUI Agent agent. We broadly categorize these strategies into two types: (1) Prompt-based Methods and (2) Training-based Methods. Prompt-based methods do not involve the training of parameters; they elicit the ability to solve agentic tasks by providing detailed instructions or demonstrations within the prompt. Training-based methods, on the other hand, involve optimizing the agent’s parameters to maximize an objective, such as pretraining, fine-tuning, or reinforcement learning.
@@ 5.1 Prompt-based Methods
One key approach is the use of dynamic action generation and accumulation. DynaSaur (Nguyen et al., 2024) enables agents to dynamically create and compose actions by generating and executing Python code via prompting. Given task instructions, the agent outputs code snippets defining new actions or reusing existing ones, effectively learning new skills and improving performance over time. Agent Q (Putta et al., 2024) and OSCAR (Wang and Liu, 2024) incorporate self-reflection and self-critique mechanisms via prompts, enabling agents to iteratively improve decision-making by identifying and rectifying errors. Auto-Intent (Kim et al., 2024) focuses on unsupervised intent discovery and utilization, extracting intents from interaction histories and incorporating them into future prompts. Other techniques include state-space exploration in LASER (Ma et al., 2023), state machine in OSCAR (Wang and Liu, 2024), expert development and multi-agent collaboration in MobileExperts (Zhang et al., 2024b), and app memory in AutoDroid (Wen et al., 2024).
@@ 5.2.1 Pre-training
Training of recent GUI agent models often involve the continual pre-training of existing vision large language models on additional large-scale datasets. This step refines the model’s general knowledge and modifies or assembles new neural network modules into the backbone, providing a stronger foundation before fine-tuning on smaller, curated datasets for GUI tasks. VisionLLM Wang et al. (2023) utilizes public datasets to integrate BERT Devlin (2018) and Deformable DETR Zhu et al. (2020) into large language models, focusing on visual question answering tasks centered on grounding and detection. SeeClick Cheng et al. (2024a) is built using continual pre-training on Qwen-VL Bai et al. (2023) with datasets incorporating OCR-based layout annotation to predict click actions. UGround Gou et al. (2024) use continual pre-training on the LLaVA-NEXT Liu et al. (2024a) model without its low-resolution image fusion module on a large dataset and synthetic data to align visual elements with HTML metadata for planning and grounding tasks.
Pre-training is also used to adapt new designs for improved computational efficiency in GUI-related tasks. CogAgent Hong et al. (2023) employs a high-resolution cross-module to process small icons and text, enhancing its efficiency for GUI tasks such as DOM element generation and action prediction. ShowUI Lin et al. (2024) built on Qwen2-VL Wang et al. (2024c) with a visual-token selection module to improve the computational efficiency for interleaved high-resolution grounding.
@@ 5.2.2 Fine-tuning
Recent work highlights reducing hallucinations and improving grounding. Falcon-UI Shen et al. (2024a) fine-tunes on large-scale instruction-free GUI data and then fine-tunes on Android and Web tasks, achieving high accuracy with fewer parameters. VGA Ziyang et al. (2024), through image-centric fine-tuning, reduces hallucinations by tightly coupling visual inputs with GUI elements, thus improving action reliability. Similarly, UI-Pro Li et al. (2024) identifies a hidden recipe for systematic fine-tuning of VLMs, scaling down model size while maintaining state-of-the-art grounding accuracy.
Other methods leverage fine-tuning to incorporate domain-specific reasoning and functionalities such as functionality-aware fine-tuning for generating human-like interactions Liu et al. (2024d), alignment strategies to handle multilingual, variable-resolution GUI inputs  Nong et al. (2024). Some methods emphasize autonomous adaptation, such as learning to execute arbitrary voice commands through trial-and-error exploration Pan et al. (2023) and learning for cross-platform GUI grounding without structured text Cheng et al. (2024a). Additionally, fine-tuning can specialize models for context-sensitive actions. Techniques proposed by Liu et al. (2023) enable context-aware text input generation, improving coverage in GUI testing scenarios. Taken together, these fine-tuning methods demonstrate how careful parameter adaptation, data scaling and multimodal alignment can collectively advance the reliability, interpretability, and performance of GUI agents.
@@ 5.2.3 Reinforcement Learning
Reinforcement learning (RL) was used in the early text-based agent WebGPT to improve information retrieval of the GPT-3 based model Nakano et al. (2021).  Liu et al. (2018) use human demonstrations to constrain the search space for RL, though using workflows as a high-level process for the model to complete without specifying the specific details. An example from  Liu et al. (2018) is for the specific process of forwarding a given email, the workflow would involve clicking forward, typing in the address, and clicking send.  Deng et al. (2023) uses RL based on human demonstrations as the reward signal. While early agents constrained the input and action spaces to only text, recent work has extended to GUI agents.
WebRL framework uses RL to generate new tasks based on previously unsuccessful attempts as a mitigation for sparse rewards Qi et al. (2024). Task success is evaluated by an LLM-based outcome reward model (ORM) and KL-divergence is used to prevent significant shifts in policies during the curriculum. AutoGLM apply online, curriculum learning, in particular to address error recovery during real-world use and to correct for stochasticity not present in simulators Liu et al. (2024c). DigiRL uses a modified advantage-weighted regression (AWR) algorithm for offline learning Peng et al. (2019), but modifies AWR for more stochastic environments by using a simple value function and curriculum learning.
@@ 6 Open Problems & Challenges
GUI Agents still struggle to accurately infer user goals across diverse applications, achieving only 51.1% accuracy on unseen websites Kim et al. (2024). Designing models that generalize effectively across varying tasks is crucial, particularly for handling contextual variations in user interactions Stefanidi et al. (2022) and predicting user behavior in complex interfaces Gao et al. (2024). Future research prospects involve leveraging robust training techniques to enable agents to adapt to new environments with minimal retraining, ultimately providing more seamless and adaptive user experiences.
@@
GUI agents often interact with sensitive data, including passwords, confidential documents, and personal credentials, raising significant privacy and security concerns He et al. (2024a); Zhang et al. (2024a). These risks are amplified when agents rely on cloud-based processing, which involves transmitting sensitive data to remote servers. Unauthorized access or incorrect information could lead to severe consequences Zhang et al. (2024c). Future research could focus on developing privacy-preserving protocols, exploring local processing alternatives, and implementing advanced authentication methods to mitigate these risks and ensure the reliability and safety of GUI agents across diverse environments.
@@
The need to handle complex interactions across diverse applications often conflicts with the demand for real-time responsiveness. Optimizing model efficiency without sacrificing accuracy remains a key hurdle, especially when deploying agents in resource-constrained environments. Challenges include reducing computational overhead, integrating hardware acceleration, and balancing trade-offs between speed and resource usage. Addressing these issues will require lightweight model designs and adaptive techniques to ensure timely, seamless interactions in dynamic GUI settings.
@@ 7 Conclusion
In this survey, we have thoroughly explored GUI Agents, examining various benchmarks, agent architectures, and training methods. Although considerable strides have been made, problems such as intent understanding, security, latency, and personalization remain critical challenges. We hope this survey will act as a valuable resource for researchers, offering structure and practical guidance in this rapidly growing and exciting field, and inspiring further inquiry into GUI Agents. We are confident that the progress in this area will mark an important milestone, benefiting humankind, significantly enhancing our daily productivity, and transforming the way we interact with computers.
@@
We recognize that some studies have explored interactions between LFM-based agents and digital systems through interfaces other than GUIs, such as Command Line Interfaces (CLI) or Application Programming Interfaces (API). However, these approaches are relatively limited in scope compared to GUI-based methods. To maintain a focused scope for our survey, we have chosen not to include them in our discussion.

### v2

@@ 4.3.2 Planning with External Knowledge
Enabling LLM-powered agents to interact with diverse applications and resources through GUIs allows them to leverage external data sources, thereby enhancing their planning capabilities. For example, Search-Agent (Koh et al., 2024b) combines LLM inference with A* search to explore and backtrack to alternative paths explicitly, AgentQ (Putta et al., 2024) combines LLM with MCTS. Toolchain (Zhuang et al., 2023) models tool planning as a tree search algorithm and incorporates A* search to adaptively retrieve the most promising tool for subsequent use based on accumulated and anticipated costs. SGC (Wu et al., 2024a) decomposes the query and performs embedding similarity match between the concatenated subquery with the current retrieved task API and each of the existing APIs, and then selects the top one from the existing neighboring APIs. Thought Propagation Retrieval (Yu et al., 2023) prompts LLMs to propose a set of analogous problems and then applies established prompting techniques, like Chain-of-Thought, to derive solutions. The aggregation module subsequently consolidates solutions from these analogous problems, enhancing the problem-solving process for the original input. Benchmarks like WebShop, Mind2Web, and WebArena Zhou et al. (2023c); Deng et al. (2023) enable agents to interact with web environments to plan and execute browsing actions for information-seeking tasks. WMA (Chae et al., 2024) utilizes world models to address the mistakes made by LLMs for long-horizon tasks.
@@ 4.4 Acting
Acting in GUI agents involves translating the agent’s reasoning and planning outputs into executable steps within the GUI environment. Unlike purely text-based or API-driven agents, GUI agents must articulate their actions at a finer granularity—often down to pixel-level coordinates—while also handling higher-level semantic actions such as typing text, scrolling, or clicking on specific elements. Several approaches have emerged:
@@ 5 GUI Agent Training Methods
This section summarizes different strategies to elicit the ability to solve agentic tasks in a GUI Agent agent. We broadly categorize these strategies into two types: (1) Prompt-based Methods and (2) Training-based Methods. Prompt-based methods do not involve the training of parameters; they elicit the ability to solve agentic tasks by providing detailed instructions within the prompt. Training-based methods, on the other hand, involve optimizing the agent’s parameters to maximize an objective, such as pretraining, fine-tuning, or reinforcement learning. We present a taxonomy of GUI agent training methods in Figure 2.
@@ 5.1 Prompt-based Methods
Agent Q (Putta et al., 2024) and OSCAR (Wang and Liu, 2024) incorporate self-reflection and self-critique mechanisms via prompts, enabling agents to iteratively improve decision-making by identifying and rectifying errors. Auto-Intent (Kim et al., 2024) focuses on unsupervised intent discovery and utilization, extracting intents from interaction histories and incorporating them into future prompts. Other techniques include state-space exploration in LASER (Ma et al., 2023), state machine in OSCAR (Wang and Liu, 2024), expert development and multi-agent collaboration in MobileExperts (Zhang et al., 2024b), and app memory in AutoDroid (Wen et al., 2024).
@@ 5.2.1 Pre-training
Training of recent GUI agent models often involve the continual pre-training of existing vision large language models on additional large-scale datasets. This step refines the model’s general knowledge and modifies or assembles new neural network modules into the backbone, providing a stronger foundation before fine-tuning on smaller, curated datasets for GUI tasks. VisionLLM Wang et al. (2023) utilizes public datasets to integrate BERT Devlin (2018) and Deformable DETR Zhu et al. (2020) into large language models, focusing on visual question answering tasks centered on grounding and detection. SeeClick Cheng et al. (2024a) is built using continual pre-training on Qwen-VL Bai et al. (2023) with datasets incorporating OCR-based layout annotation to predict click actions. UGround Gou et al. (2024) uses continual pre-training on the LLaVA-NEXT Liu et al. (2024a) model without its low-resolution image fusion module on a large dataset and synthetic data to align visual elements with HTML metadata for planning and grounding tasks.
Pre-training is also used to adapt new designs for improved computational efficiency in GUI-related tasks. CogAgent Hong et al. (2023) employs a high-resolution cross-module to process small icons and text, enhancing its efficiency for GUI tasks such as DOM element generation and action prediction. ShowUI Lin et al. (2024) builds on Qwen2-VL Wang et al. (2024c) with a visual-token selection module to improve the computational efficiency for interleaved high-resolution grounding.
@@ 5.2.2 Fine-tuning
Recent work highlights reducing hallucinations and improving grounding. Falcon-UI Shen et al. (2024a) fine-tunes on large-scale instruction-free GUI data and then fine-tunes on Android and Web tasks, achieving high accuracy with fewer parameters. VGA Ziyang et al. (2024), through image-centric fine-tuning, reduces hallucinations by tightly coupling visual inputs with GUI elements, thus improving action reliability. Similarly, UI-Pro Li et al. (2024) identifies a recipe for fine-tuning of VLMs, reducing model size while maintaining state-of-the-art grounding accuracy.
Other methods leverage fine-tuning to incorporate domain-specific reasoning and functionalities, such as functionality-aware fine-tuning for generating human-like interactions Liu et al. (2024d) and alignment strategies to handle multilingual, variable-resolution GUI inputs  Nong et al. (2024). Some methods emphasize autonomous adaptation, such as learning to execute arbitrary voice commands through trial-and-error exploration Pan et al. (2023) and learning for cross-platform GUI grounding without structured text Cheng et al. (2024a). Additionally, fine-tuning can specialize models for context-sensitive actions. Techniques proposed by Liu et al. (2023) enable context-aware text input generation, improving coverage in GUI testing scenarios. Taken together, these fine-tuning methods demonstrate how careful parameter adaptation, data scaling and multimodal alignment can collectively advance the reliability, interpretability, and performance of GUI agents.
@@ 5.2.3 Reinforcement Learning
Reinforcement learning was used in the early text-based agent WebGPT to improve information retrieval of the GPT-3 based model Nakano et al. (2021).  Liu et al. (2018) use human demonstrations to constrain the search space for RL, through using workflows as a high-level process for the model to complete without specifying the specific details. An example from  Liu et al. (2018) is for the specific process of forwarding a given email, the workflow would involve clicking forward, typing in the address, and clicking send.  Deng et al. (2023) use RL based on human demonstrations as the reward signal. While early agents constrained the input and action spaces to only text, recent work has extended to GUI agents.
WebRL framework uses RL to generate new tasks based on previously unsuccessful attempts as a mitigation for sparse rewards Qi et al. (2024). Task success is evaluated by an LLM-based outcome reward model (ORM) and KL-divergence is used to prevent significant shifts in policies during curriculum learning. AutoGLM applies online, curriculum learning, in particular to address error recovery during real-world use and to correct for stochasticity not present in simulators Liu et al. (2024c). DigiRL uses a modified advantage-weighted regression (AWR) algorithm for offline learning Peng et al. (2019), but modifies AWR for more stochastic environments by using a simple value function and curriculum learning.
@@ 6 Open Problems & Challenges
GUI Agents still struggle to accurately infer user goals across diverse applications, achieving only 51.1% accuracy on unseen websites Kim et al. (2024). Designing models that generalize effectively across varying tasks is crucial, particularly for handling contextual variations in user interactions Stefanidi et al. (2022) and predicting user behavior in complex interfaces Gao et al. (2024). A prospective future research direction is to leverage robust training techniques to enable agents to adapt to new environments with minimal retraining, ultimately providing more seamless and adaptive user experiences. Other promising directions could include training on diverse user interaction datasets and incorporating context-aware learning techniques that utilize historical user actions to better predict intent.
@@
GUI agents frequently interact with sensitive data such as passwords, confidential documents, and personal credentials, raising serious privacy and security concerns He et al. (2024a); Zhang et al. (2024a). These risks are further amplified when agents rely on cloud-based processing, which involves transmitting sensitive information to remote servers. Unauthorized access or incorrect actions could result in severe consequences Zhang et al. (2024c). Future research could focus on developing privacy-preserving protocols, such as homomorphic encryption or differential privacy, to ensure data remains secure during both inference and storage. Additional directions may include exploring local processing alternatives and implementing advanced authentication mechanisms to enhance the reliability and safety of GUI agents across diverse environments.
@@
The need to manage complex interactions across diverse applications often conflicts with the requirement for real-time responsiveness. Optimizing model efficiency without compromising accuracy remains a key challenge, particularly when deploying agents in resource-constrained environments. Key issues include minimizing computational overhead, leveraging hardware acceleration, and balancing trade-offs between speed and resource usage. Addressing these challenges calls for lightweight model architectures and adaptive techniques that enable timely, seamless interactions in dynamic GUI settings. Future research could investigate hardware-aware optimization methods, such as quantization and pruning, or efficient decoding strategies like predictive sampling and multi-token prediction, which can significantly reduce latency while preserving system accuracy.
Personalization.
is a pivotal aspect in the development of GUI agents, aiming to tailor interactions to individual user preferences and behaviors, thereby enhancing satisfaction and efficiency. Recent work Berkovitch et al. (2024) introduced a method for identifying user goals from UI trajectories, enabling agents to infer intentions and proactively assist users based on their interactions with the interface. Future research could explore more sophisticated models that incorporate user feedback to refine personalization strategies, while ensuring trust and compliance with data protection regulations. Additional directions include implementing explicit feedback mechanisms (e.g., thumbs-up/thumbs-down ratings) and developing robust user profiling techniques that integrate behavioral and contextual data to enable more meaningful and adaptive personalization.
@@ 7 Conclusion
In this survey, we have thoroughly explored GUI Agents, examining various benchmarks, agent architectures, and training methods. Although considerable strides have been made, problems such as intent understanding, security, latency, and personalization remain critical challenges. We hope that this survey is a valuable resource for researchers, offering structure and practical guidance in this rapidly growing and exciting field, and inspiring more work on GUI Agents. The progress in this area has already benefited mankind, enhancing our daily productivity and transforming the way we interact with computers.
@@
We recognize that some studies have explored interactions between LFM-based agents and digital systems through interfaces other than GUIs, such as Command Line Interfaces (CLI) (Nguyen et al., 2024) or Application Programming Interfaces (API). (Song et al., 2025) However, these approaches are relatively limited in scope compared to GUI-based methods. To maintain a focused scope for our survey, we have chosen not to include them in our discussion.